OpenStudy (anonymous):

Help please? Solve 2x2 - 8x = -7

OpenStudy (anonymous):

So first move all terms to one side

OpenStudy (anonymous):

Yes, like i did with the last equation, do I have to put this one into the quadratic formula or no?

OpenStudy (anonymous):

Yes you do

OpenStudy (anonymous):

I did that but I think i did it wrong because I didn't get any of the choices

OpenStudy (anonymous):

negative 2 plus or minus square root of 2 negative 2 plus or minus 2 square root of 2 quantity of 2 plus or minus square root of 2 all over 2 2 plus or minus square root of 2 end root over 2

OpenStudy (phi):

start with \[ 2x^2 - 8x = -7 \] add +7 to both sides, and you get \[ 2x^2 - 8x +7=0 \] what are "a", "b" and "c" for this equation? in other words, match up your equation with \[ ax^2 +bx+c= 0 \]

OpenStudy (phi):

If you want more info, maybe this helps https://www.khanacademy.org/math/algebra/quadratics/quadratic-formula/v/using-the-quadratic-formula

OpenStudy (anonymous):

I know how to do the beginning to this equation but once I get to the solving it in the quadratic formula I get messed up, @phi

OpenStudy (anonymous):

@kidrah69 help?

OpenStudy (kidrah69):

Have you figured out what your a b and c value is?

OpenStudy (anonymous):

a, 2 b, -8 c, -7

OpenStudy (kidrah69):

\[\frac{ x = -b ± \sqrt(b^2 - 4ac) }{ 2a }\] good now plug it in

OpenStudy (anonymous):

When I plug all that in i get -8 plus or minus the srqrt of -8^2 - 4(2)(-7) all over 2(2)

OpenStudy (kidrah69):

\[\sqrt{(-8)^2 - 4(2)(-7)}\] solve

OpenStudy (anonymous):

64-56 =8/4

OpenStudy (kidrah69):

no just solve whats in the square root dont square root it :P

OpenStudy (anonymous):

what do you mean? if i take the -8^2 it will be 64 and then 4 times the 2 and the -7 would be 56 which equals 8?

OpenStudy (kidrah69):

hmmm we are not doing something right.....

OpenStudy (anonymous):

I'm confused?

OpenStudy (kidrah69):

Well one problem it instead of a -7 there should be just a 7

OpenStudy (anonymous):

ahh you're right cause we added it to the other side

OpenStudy (kidrah69):

We still get 8 though :/

OpenStudy (anonymous):

yeah that's my problem, and the answer choices don't add up with my math

OpenStudy (anonymous):

negative 2 plus or minus square root of 2 negative 2 plus or minus 2 square root of 2 quantity of 2 plus or minus square root of 2 all over 2 2 plus or minus square root of 2 end root over 2

OpenStudy (kidrah69):

2x2 - 8x = -7 Quadratic Formula: \(x=\dfrac{8\pm\sqrt{4^2-4(2)(7)}}{2(2)}\) \(x=\dfrac{8\pm\sqrt{16-56}}{4}\) \(x=\dfrac{8\pm\sqrt{40}}{4}\)

OpenStudy (anonymous):

But that still isn't any of the choices. It's okay don't stress this one, could you help me with something else?

OpenStudy (kidrah69):

Wat is the topic of the other one?

OpenStudy (anonymous):

Same thing, I just want to know sometimes the answer have the same thing but one will be positive and the other answer negative (still dealing with quadratic formula) how do I know if its pos or neg?

OpenStudy (kidrah69):

Is this solving the formula? or do u mean with the abc values

OpenStudy (anonymous):

I think part of the formula.. look quantity of negative 9 plus or minus 3 square root of 5 all over 2 quantity of 9 plus or minus 3 square root of 5 all over 2 see how their the same answer but one is a negative 9 and the other a positive? How do I know which one is correct?

OpenStudy (anonymous):

@kidrah69, do you get what I'm saying?

OpenStudy (kidrah69):

Well when you do the work it before the squrt is has -9 it will be a negative number if not assume its postivie

OpenStudy (anonymous):

Okay, thanks for you help! I appreciate it :-)

OpenStudy (anonymous):

thank you soo mcuh!

OpenStudy (kidrah69):

:)

OpenStudy (phi):

\[ 2x^2 - 8x +7=0 \] a=2 , b= -8, c=7 now do things slowly and carefully \[ b^2 -4 ac= (-8)^2 - 4\cdot 2 \cdot 7= 64 - 56 = 8\] \[ \sqrt{ b^2 -4 ac}= \sqrt{8} \\= \sqrt{4\cdot 2} \\= \sqrt{4} \sqrt{2} \\= 2 \sqrt{2}\] so we the answers are \[ \frac{-b \pm \sqrt{ b^2 -4 ac}}{2a} \\= \frac{-(-8) \pm 2 \sqrt{2}}{2\cdot 2} \] you can simplify that to \[ \frac{8 \pm 2 \sqrt{2}}{2\cdot 2} = 2 \pm \frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

Thank you so much @phi! :)